If a sequence is defined recursively by k(0)=2 and k(x+1) = 3k(x)-5 for x is greater than or equal to 0, then k(3) is equal to what value?
Answer by stanbon(75887)

If a sequence is defined recursively by k(0)=2 and k(x+1) = 3k(x)-5 for x is greater than or equal to 0, then k(3) is equal to what value?
----
k(0)= 2
k(x+1) = 3k(x)-5
-----------------
k(1) = 3*2-5 = 1
------------------
k(2) = 3*1-5 = -2
-------
k(3) = 3*(-2)-5 = -11
=============
Cheers,
Stan H.
================